dy If y(x) is the solution of dr 22 +1 and y(0) = 2 then y(3) = 33 OA +2 OB. +2V2 ос 0 OD +4 OE +4V2
Answer #1

Given question solved. of Y (²) is the solution he x+1 y? and y (o)=2 theny (3) dy x²+1 da =7% dy y? (x²41) dx taking integration both sidsy Ju dy -4 十 [4( + +) Now, 九 ||| ti +3+4 3 Yy | 4(1+2+1) - [1(+0] 4 M *[Axle] 4x16 - 64 *2 > 十 [26] ML 17 } + [2] YL 十 +25 • + 2] so =